The table below shows Louisa’s preferences for the number of books she buys per month, holding...

  1. The table below shows Louisa’s preferences for the number of books she buys per month, holding the prices of all goods constant.

Income (thousands of $)

Optimal number of books

Income elasticity of demand for books

10

3

20

5

30

10

40

15

50

15

  1. Complete the income elasticity of demand column for the table above (using the midpoint method). Are books ever an inferior good for Louisa? At what level(s) of income are books a luxury?
  2. Draw a smooth approximation of Louisa’s Engel curve for books.
Answer #1

optimal numba I of Income (thousands of demand Income elasticity i for books books 10 xx xx 10:75 10 O 1.67 idos; ensino sulfor the income range of $20-$30 (m) i ind pucentage change in income 130-20) *100 = 40 130720) 2 Quantity demanded for booksfor the income range of $30-$40% pucentage change in income (h) - 10 28.57 = 40-30) 140+30) 2 for books (ad): and pucentage cAs value of in come elasticity of demand is positive at all price ranges, hence books are neun an inferior good to Louisa. NoIncome (thousands of Engel cmue for books . 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 is optimal » bu books CS Scanned with CamScan